OpenStudy (anonymous):

What is the volume of a square pyramid with base edges of 60 inches and a height of 10 inches? 600 in3 6,000 in3 12,000 in3 36,000 in3

OpenStudy (anonymous):

i'm assuming it's 12,000 but i wasn't fully sure

OpenStudy (anonymous):

Do you know the formula?

OpenStudy (anonymous):

you are correct though

OpenStudy (anonymous):

First you'd find the base. Then the volume . Volume: 1/3 Bh Base:S squared

OpenStudy (anonymous):

oh ok.

OpenStudy (anonymous):

thats because the formula is v = 1/3b^2h v = 1/3(60^2)(10) v =1/3(3600)(10) v = 1/3(36000) v = 12000